# Astrocyte Regulated Neuromorphic Central Pattern Generator Control of Legged Robotic Locomotion

### v2

@@ V-D Power Consumption Benefits
Recent literature on conventional deep learning approaches for quadruped robot gait learning primarily involve reinforcement learning where the action at each control time step is output from a policy network. For example, [9] introduces an energy-related reward function in a reinforcement learning setting for quadruped (Unitree A1 robot) locomotion. The policy network is a 3-layer architecture where each hidden layer contains 128 neurons and the output layer contains 12 neurons providing the target angular positions of the 12 joints in the limbs. The input to the policy network is a 42-dimensional state vector containing the current angular position and velocity of the joints in the limbs, the roll and pitch of the torso, the foot contact indicators and the action produced in the previous control time step. Therefore, the power consumption of the control system can be regarded as the product of the energy consumption of a single feed-forward inference computation and the control frequency. For the feed-forward inference, the number of multiplication operations of a single layer $j$ with the bias of an artificial neural network is:
@@
where $[(d_{i}-1)+1]$ represents the number of additions including bias for a single output neuron. Therefore, the power consumption of the policy network based control system can be calculated as,
@@
where, $f_{\mathrm{control}}=100$Hz is the control frequency [9], $\epsilon_{\mathrm{MULT}}$ and $\epsilon_{\mathrm{ADD}}$ are the energy consumption of a single multiplication and a single addition operation. In this paper, we consider the multiplication and addition operations are implemented in $45$nm CMOS process [16], where $\epsilon_{\mathrm{MULT}}=3.7\mathrm{pJ}$ and $\epsilon_{\mathrm{ADD}}=0.9\mathrm{pJ}$. Consequently,
@@
The SNN-based CPG proposed in this paper operates by means of transmitting and receiving spikes. Instead of the multiply and addition operation that takes place for each synaptic connection in a non-spiking network, only an ADD operation takes place per synaptic operation in an SNN upon the receipt of an incoming spike [37]. The energy consumption of the SNN-based control system therefore can be calculated by counting the total number of accumulation operations. The discrete simulation of an SNN implies that the simulation frequency does not impact the firing frequencies of the spiking neurons. Thus the power consumption of the SNN-based CPG is independent of the SNN simulation frequency, excluding the sensory input portion in this paper, which will be discussed in details below. As elaborated in the Methods section, the neurons composing the CPG can be categorized into motor neurons and inhibitory interneurons including the V1/V2b interneurons maintaining the reciprocal inhibition and the IINs between the thigh and the calf muscles in the same limb. These two types of inhibitory interneurons both individually inhibit 20 corresponding motor neurons, and hence they can be treated uniformly: one spike sent from an inhibitory interneuron results in 20 accumulation operations. One motor neuron in the calf flexor and extensor motor neuron pool excites 19 other motor neurons in the same motor neuron pool and excites a V1/V2b interneuron, which implies that one spike from a calf muscle motor neuron also leads to 20 accumulation operations. The case is slightly different for the thigh motor neurons. In addition to exciting 19 motor neurons in the same motor neuron pool and a V1/V2b interneuron, one thigh motor neuron also excites an IIN to inhibit the antagonistic muscle in the calf. Additionally, one thigh motor neuron in one limb sends 60 axons in total to the thigh motor neuron pools in all the other limbs. Thus one spike from a thigh muscle motor neuron results in 81 accumulation operations. Also, when a thigh joint is approaching the limit positions, the 20 corresponding thigh muscle motor neurons receive limit position inhibitory spikes. The limit position spiking frequency here is the simulation frequency of the SNN, which is 1000Hz. Increasing the SNN simulation frequency will not influence the strength of limit position inhibition onto the motor neurons, but it will slightly raise the total power consumption. The total power consumption of the SNN-based controller is:
@@
where, $f_{\mathrm{all\_inhi}}$, $f_{\mathrm{all\_calf}}$, $f_{\mathrm{all\_thigh}}$ and $f_{\mathrm{all\_limit\_posistion}}$ are the total firing frequencies of all the inhibitory interneurons, calf muscle motor neurons, thigh muscle motor neurons and limit position inhibition. The total firing frequencies of the neurons are estimated by taking the mean of session-wise average firing frequencies of the neurons. Specifically, the CPG runs for a session whose maximum session length is 10s to produce one session-wise average firing frequency. The estimated total firing frequency is the average of 10 session-wise average firing frequencies. The estimated average firing frequencies are $f_{\mathrm{all\_inhi}}=6.69\times 10^{2}\mathrm{Hz}$, $f_{\mathrm{all\_calf}}=4.34\times 10^{3}\mathrm{Hz}$, $f_{\mathrm{all\_thigh}}=4.52\times 10^{3}\mathrm{Hz}$ and $f_{\mathrm{all\_limit\_posisition}}=2.22\times 10^{3}\mathrm{Hz}$. Consequently, $P_{\mathrm{SNN-CPG}}=4.60\times 10^{-7}\,\,\mathrm{Watt}$, which is 23.3×23.3$\times$23.3 × power efficient than the reinforcement learning based approach.
@@ VI Discussion
In summary, the work paves the way for energy-efficient neuromorphic CPG control of legged robotic locomotion tasks in edge robotics applications and underscores the critical homeostatic role of astrocytes in the gait emergence process.
Future work should consider further strengthening of the neuroscience connection in CPG modelling to develop adaptive and energy-efficient robotic locomotion. For instance, a more bio-plausible CPG modelling can consider the torque applied to a joint to have a complex relationship with the background stimulation. Multiple types of interneurons in the spinal cord, in fact, are involved in the neural circuit at different locomotion speeds [20], which could be a topic of further research. Descending command signals from the brainstem and sensory inputs are critical for animal locomotion. Also, the muscles in the torso are indispensable in multiple types of action. At first glance of animal locomotion, it seems that the legs are contributing most to the emergence of gaits. However, the limb muscles are only a fraction of the actuators of animal movement. The feed-forward plus feedback organization of all the muscles and the sensory apparatus as an entity is the true nature of animal locomotion and underlie the locomotion efficiency of biological systems.
@@ Acknowledgments
A single limb of the A1 robot contains three joints which are called hip, thigh and calf, which can be noticed in Figure 1d. The torque outputs from the central pattern generator are directly connected to the thigh and calf joint motor actuators. The terrain is flat ground. The robot’s initial heading direction is the positive $x$ direction. The thigh and calf joint limit positions are modified from the manufacturer’s setting for better gait formation, and the hip joint limit position is not changed. The position range for the front and rear thigh joints are correspondingly $[\mathrm{pos^{limit,-}_{thigh,front}},\mathrm{pos^{limit,+}_{thigh,front}}]=[0.6,1.4]$ and $[\mathrm{pos^{limit,-}_{thigh,rear}},\mathrm{pos^{limit,+}_{thigh,rear}}]=[0.7,1.5]$. The position range for the calf joint is $[\mathrm{pos^{limit,-}_{calf}},\mathrm{pos^{limit,+}_{calf}}]=[-1.6,-1.0]$, all position values are in unit rad. As described in the Locomotion Unit section in the main text, the inhibition region for the thigh joint is the region within 0.05 from the limit positions. The “frictionloss” attribute of the joints is also increased to prevent fast swing angular velocity. The “frictionloss” parameter for hip and calf joints is set to 10 while that for the thigh joint is set to 25. The hip joints allow the limbs to move medially or laterally, therefore they do not significantly contribute to the forward driving force onto the torso. Therefore a simple PI controller is designed to stabilize the hip joints at a constant position, where $K_{\mathrm{P}}=30$ and $K_{\mathrm{I}}=10$. The designed stable position of the hip joints is laterally $0.1$, i.e. $0.1$ rad away from the center line of the torso. The expression of the hip joint torque as a function of the current and accumulative position error therefore is
[…]
@@
Before the onset of each training session, $t_{\mathrm{learning\_start}}$ and $\mathrm{Progress}(t)$ are calculated. The hip PI controller, motor traces, MuJoCo system status and all the SNN neuron status are reset. The astrocyte status is not reset before the beginning of a new session since the $\mathrm{Ca}^{2+}$ dynamics is a slow fluctuating process that will not have abrupt change because of the change of robot status and is a reflection of the average motor neuron activities over several seconds. The thigh and calf joint positions are reset to a weighted sum of the limit positions of the joints. Accurately:
@@
Parameter
Description
Value
@@
$w_{0\mathrm{dist}}$
Intra-pool zero distance synapse strength
4
$c_{\mathrm{dist}}$
Synapse strength spatial decay coefficient
0.3
$w_{\mathrm{V1/V2b}\xrightarrow{}\mathrm{motor}}$
Uniform synapse strength from V1/V2b to motor neuron
-50
$w_{\mathrm{motor}\xrightarrow{}\mathrm{V1/V2b}}$
Uniform synapse strength from motor neuron to V1/V2b
2
@@
$\mathbf{Size}_{\mathrm{pool}}$
Number of motor neurons in a motor neuron pool
$20$
$v_{\mathrm{motor\_rest}}$
Motor neuron resting and reseting potential
00 mV
$v_{\mathrm{motor\_th}}$
Motor neuron threshold potential
$10$ mV
$\tau_{\mathrm{motor}}$
Motor neuron membrane potential decay time constant
$9$ ms
$\mathrm{refractory}_{\mathrm{motor}}$
Motor neuron refractory period
$5$ ms
$S_{\mathrm{motor\_spike}}$
Motor neuron stochastic firing transition width
$0.2$
$r_{\mathrm{Ca^{2+}}}$
$[\mathrm{Ca^{2+}}]$ unit increment per action potential
$1$
$\tau_{\mathrm{Ca^{2+}}}$
$[\mathrm{Ca^{2+}}]$ decay time constant
$250$ ms
$\mathrm{Thres}_{\mathrm{Ca}^{2+}}$
$[\mathrm{Ca^{2+}}]$ threshold for opening the $\mathrm{K^{+}}$ channel
$10$
$S_{\mathrm{K\_chan}}$
$[\mathrm{Ca^{2+}}]$ sensitivity in $\mathrm{K^{+}}$ channel
$10$
$C_{\mathrm{K\_chan}}$
$\mathrm{K^{+}}$ channel induced potential drop rate
$8000$ mV/s
$C_{\mathrm{limit\_position}}$
Limit position inhibitory potential drop rate
$400$ mV/s
$I_{\mathrm{background},0}$
Zero-velocity background stimulation strength
$1380$ mV/s
$k_{\mathrm{background},v}$
Velocity-dependent background stimulation increment coefficient
[…]
$\mathrm{A}_{\mathrm{random}}$
Randomness amplitude of background stimulation
0.5
@@
$v_{\mathrm{V1/V2b\_rest}}$
V1/V2b resting and reseting potential
00 mV
$v_{\mathrm{V1/V2b\_th}}$
V1/V2b threshold potential
$10$ mV
$\tau_{\mathrm{V1/V2b}}$
V1/V2b membrane potential decay time constant
$9$ ms
$\mathrm{refractory}_{\mathrm{V1/V2b}}$
V1/V2b refractory period
$3$ ms
$S_{\mathrm{V1/V2b\_spike}}$
V1/V2b stochastic firing transition width
$0.2$
@@
$\tau_{h}$
Motor torque trace decay time constant
$0.1$ s
$I_{\mathrm{thigh}}$
Torque increment for a single spike at thigh joint
$0.7$ $\mathrm{N}\cdot\mathrm{m}$
$I_{\mathrm{calf}}$
Torque increment for a single spike at calf joint
$1.1$ $\mathrm{N}\cdot\mathrm{m}$
$\mathrm{pos^{*}_{hip}}$
Hip joints’ target position
lateral $0.1$ rad
$K_{\mathrm{P}}$
Hip controller P coefficient
$30$
$K_{\mathrm{I}}$
Hip controller I coefficient
$10$
@@
Parameter
Description
Value
@@
$[w_{\mathrm{min}},w_{\mathrm{max}}]$
Weight constraint factor
$[-0.05,0.05]$
$\tau_{\mathrm{trace}}$
Motor neuron spike trace decay time constant
$0.01$ s
$\tau_{\mathrm{STDP}}$
STDP signal decay time constant
$2$ s
$\eta_{\mathrm{negative\_relative}}$
Relative negative learning rate in STDP signal
$0.3$
$\eta$
Learning rate
$5\times 10^{-10}$
$c_{\mathrm{average}}$
Averaged reward coefficient
$0.5$
$\alpha_{\mathrm{vel}_{x}}$
$x$-directional velocity reward coefficient
$1$
$\alpha_{\omega_{\mathrm{roll}}}$
Roll angular velocity reward coefficient
$-0.1$
$\alpha_{\omega_{\mathrm{pitch}}}$
Pitch angular velocity reward coefficient
$-0.1$
$\alpha_{\omega_{\mathrm{yaw}}}$
Yaw angular velocity reward coefficient
$-0.1$
@@
$r_{\mathrm{AG}}$
2-AG production rate
$1\times 10^{-3}$
$\tau_{\mathrm{AG}}$
[2-AG] decay time constant
$1$ s
$\mathrm{Thres\_Ca}_{\mathrm{ADO}}$
$[\mathrm{Ca}^{2+}]$ threshold for adenosine release
$0.3$
$r_{\mathrm{ADO}}$
Adenosine production rate
$0.01$
$\tau_{\mathrm{AG}}$
Adenosine decay time constant
$1$ s
$\mathrm{refractory}_{\mathrm{ADO}}$
Adenosine release refractory period
$0.3$ s
$\eta_{\mathrm{ADO}}$
Adenosine inhibition efficacy
$1.8\times 10^{-5}$
@@
$f_{\mathrm{simulation}}$
MuJoCo simulation frequency
$1000$ Hz
$L_{\mathrm{max}}$
Maximum session length
$10$ s
$L_{\mathrm{non\_alive\_threshold}}$
Non-alive counter threshold
$0.5$ s
$\mathrm{Thres}_{z\_\mathrm{ALIVE}}$
Alive torso vertical-ness threshold
$0.5$
$[\mathrm{pos^{limit,-}_{thigh,front}},\mathrm{pos^{limit,+}_{thigh,front}}]$
Front thigh joint limit position
$[0.6,1.4]$ rad
$[\mathrm{pos^{limit,-}_{thigh,rear}},\mathrm{pos^{limit,+}_{thigh,rear}}]$
Rear thigh joint limit position
$[0.7,1.5]$ rad
$[\mathrm{pos^{limit,-}_{calf}},\mathrm{pos^{limit,+}_{calf}}]$
Calf joint limit position
$[-1.6,-1.0]$ rad
$\mathrm{Z_{torso,initial}}$
Initial torso altitude
$0.35$ m
$\mathrm{frictionloss_{thigh}}$
Thigh joint frictionloss
$25$
$\mathrm{frictionloss_{calf}}$
Calf joint frictionloss
$10$
$\mathrm{frictionloss_{hip}}$
Hip joint frictionloss
$10$

### v3

@@ V-D Power Consumption Benefits
Recent literature on conventional deep learning approaches for quadruped robot gait learning primarily involves reinforcement learning, where the action at each control time step is output from a policy network. For example, [20] introduces an energy-related reward function in a reinforcement learning setting for quadruped (Unitree A1 robot) locomotion. The policy network is a 3-layer architecture where each hidden layer contains 128 neurons and the output layer contains 12 neurons, providing the target angular positions of the 12 joints in the limbs. The input to the policy network is a 42-dimensional state vector containing the current angular position and velocity of the joints in the limbs, the roll and pitch of the torso, the foot contact indicators and the action produced in the previous control time step. Therefore, the power consumption of the control system can be regarded as the product of the energy consumption of a single feed-forward inference computation and the control frequency. For the feed-forward inference, the number of multiplication operations of a single layer $j$ with the bias of an artificial neural network is:
@@
where $(d_{i}-1)+1]$ represents the number of additions, including bias, for a single output neuron. Therefore, the power consumption of the policy network based control system can be calculated as,
@@
where, $f_{\mathrm{control}}=100$Hz is the control frequency [20], $\epsilon_{\mathrm{MULT}}$ and $\epsilon_{\mathrm{ADD}}$ are the energy consumption of a single multiplication and a single addition operation. In this paper, we consider the multiplication and addition operations are implemented in 4545nm CMOS process [51], where $\epsilon_{\mathrm{MULT}}=3.7\mathrm{pJ}$ and $\epsilon_{\mathrm{ADD}}=0.9\mathrm{pJ}$. Consequently,
@@
The SNN-based CPG proposed in this paper operates by means of transmitting and receiving spikes. Instead of the multiplication and addition operation that takes place for each synaptic connection in a non-spiking network, only an ADD operation takes place per synaptic operation in an SNN upon the receipt of an incoming spike [52]. The energy consumption of the SNN-based control system therefore can be calculated by counting the total number of accumulation operations. The discrete simulation of an SNN implies that the simulation frequency does not impact the firing frequencies of the spiking neurons. Thus, the power consumption of the SNN-based CPG is independent of the SNN simulation frequency, excluding the sensory input portion in this paper, which will be discussed in detail below. As elaborated in Section III, the neurons composing the CPG can be categorized into motor neurons and inhibitory interneurons, including the V1/V2b interneurons maintaining the reciprocal inhibition and the IINs between the thigh and the calf muscles in the same limb. These two types of inhibitory interneurons both individually inhibit 20 corresponding motor neurons, and hence they can be treated uniformly: one spike sent from an inhibitory interneuron results in 20 accumulation operations. One motor neuron in the calf flexor and extensor motor neuron pool excites 19 other motor neurons in the same motor neuron pool and excites a V1/V2b interneuron, which implies that one spike from a calf muscle motor neuron also leads to 20 accumulation operations. The case is slightly different for the thigh motor neurons. In addition to exciting 19 motor neurons in the same motor neuron pool and a V1/V2b interneuron, one thigh motor neuron also excites an IIN to inhibit the antagonistic muscle in the calf. Additionally, one thigh motor neuron in one limb sends 60 axons in total to the thigh motor neuron pools in all the other limbs. Thus, one spike from a thigh muscle motor neuron results in 81 accumulation operations. Also, when a thigh joint is approaching the limit positions, the 20 corresponding thigh muscle motor neurons receive limit position inhibitory spikes. The limit position spiking frequency here is the simulation frequency of the SNN, which is 1000Hz. Increasing the SNN simulation frequency will not influence the strength of limit position inhibition onto the motor neurons, but it will slightly raise the total power consumption. The total power consumption of the SNN-based controller is:
@@
where, $f_{\mathrm{all\_inhi}}$, $f_{\mathrm{all\_calf}}$, $f_{\mathrm{all\_thigh}}$ and $f_{\mathrm{all\_limit\_posistion}}$ are the total firing frequencies of all the inhibitory interneurons, calf muscle motor neurons, thigh muscle motor neurons and limit position inhibition. The total firing frequencies of the neurons are estimated by taking the mean of the session-wise average firing frequencies of the neurons. Specifically, the CPG runs for a session whose maximum session length is 10s to produce one session-wise average firing frequency. The estimated total firing frequency is the average of 10 session-wise average firing frequencies. The estimated average firing frequencies are $f_{\mathrm{all\_inhi}}=6.69\times 10^{2}\mathrm{Hz}$, $f_{\mathrm{all\_calf}}=4.34\times 10^{3}\mathrm{Hz}$, $f_{\mathrm{all\_thigh}}=4.52\times 10^{3}\mathrm{Hz}$ and $f_{\mathrm{all\_limit\_posisition}}=2.22\times 10^{3}\mathrm{Hz}$. Consequently, $P_{\mathrm{SNN-CPG}}=4.60\times 10^{-7}\,\,\mathrm{Watt}$, which is $23.3\times$ power efficient than the reinforcement learning based approach.
@@ VI Discussion
In summary, the work proposes a bio-inspired, SNN-based control architecture for quadruped robotic systems regulated by the homeostasis property of astrocytes. The neuromorphic CPG reveals the high energy efficiency of a spike-based locomotion control system. Also, the inhibitory astrocytic regulation causing the gait emergence underscores the significant role of modeling astrocytes for neuromorphic robotic control deployed in edge robotic applications.
Future work should consider further strengthening of the neuroscience connection in CPG modeling to develop adaptive and energy-efficient robotic locomotion. For instance, a more bio-plausible CPG modeling can consider the torque applied to a joint to have a complex relationship with the background stimulation. Multiple types of interneurons in the spinal cord, in fact, are involved in the neural circuit at different locomotion speeds [41], which could be a topic of further research. Descending command signals from the brainstem and sensory inputs are critical for animal locomotion. Also, the muscles in the torso are indispensable in multiple types of action. At first glance of animal locomotion, it seems that the legs are contributing most to the emergence of gaits. However, the limb muscles are only a fraction of the actuators of animal movement. The feed-forward plus feedback organization of all the muscles and the sensory apparatus as an entity is the true nature of animal locomotion and underlies the locomotion efficiency of biological systems.
@@ Acknowledgments
Research was sponsored primarily by the Army Research Office and was accomplished under Grant Number W911NF-24-1-0127. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was also supported partially by the National Science Foundation under award No. EFRI BRAID #2318101.
@@
A single limb of the A1 robot contains three joints which are called hip, thigh and calf, which can be noticed in Figure 1d. The torque outputs from the central pattern generator are directly connected to the thigh and calf joint motor actuators. The terrain is flat ground. The robot’s initial heading direction is the positive $x$ direction. The thigh and calf joint limit positions are modified from the manufacturer’s setting for better gait formation, and the hip joint limit position is not changed. The position range for the front and rear thigh joints are correspondingly $\mathrm{pos^{limit,-}_{thigh,front}},\mathrm{pos^{limit,+}_{thigh,front}}]=[0.6,1.4]$ and $\mathrm{pos^{limit,-}_{thigh,rear}},\mathrm{pos^{limit,+}_{thigh,rear}}]=[0.7,1.5]$. The position range for the calf joint is $\mathrm{pos^{limit,-}_{calf}},\mathrm{pos^{limit,+}_{calf}}]=[-1.6,-1.0]$, all position values are in unit rad. As described in the Locomotion Unit section in the main text, the inhibition region for the thigh joint is the region within 0.05 from the limit positions. The “frictionloss” attribute of the joints is also increased to prevent fast swing angular velocity. The “frictionloss” parameter for hip and calf joints is set to 10 while that for the thigh joint is set to 25. The hip joints allow the limbs to move medially or laterally. Therefore, they do not significantly contribute to the forward driving force onto the torso. Therefore, a simple PI controller is designed to stabilize the hip joints at a constant position, where $K_{\mathrm{P}}=30$ and $K_{\mathrm{I}}=10$. The designed stable position of the hip joints is laterally $0.1$, i.e. $0.1$ rad away from the center line of the torso. The expression of the hip joint torque as a function of the current and cumulative position error therefore is
$\begin{split}\mathrm{torque_{hip}}(t)=&K_{\mathrm{P}}\cdot[\mathrm{pos^{*}_{hip}}-\mathrm{pos_{hip}}(t)]\\ +&K_{\mathrm{I}}\cdot\int_{0}^{t}[\mathrm{pos^{*}_{hip}}-\mathrm{pos_{hip}}(t)]dt\end{split}$ (22)
@@
Before the onset of each training session, $t_{\mathrm{learning\_start}}$ and $\mathrm{Progress}(t)$ are calculated. The hip PI controller, motor traces, MuJoCo system status and all the SNN neuron status are reset. The astrocyte status is not reset before the beginning of a new session since the $\mathrm{Ca}^{2+}$ dynamics is a slow fluctuating process that will not have an abrupt change because of the change of robot status and is a reflection of the average motor neuron activities over several seconds. The thigh and calf joint positions are reset to a weighted sum of the limit positions of the joints. Accurately:
@@
Computational Implementation of the Dynamical System
The STDP signal, astrocyte calcium concentration, neuron membrane potential, and the other system dynamics mentioned in this work are mostly described through a set of differential equations which have a common form. The finite difference method is adopted to implement the evolution of such dynamical systems. Figure 4 illustrates the derivation of an implementable model state updating method. For uniformity, the time constants in all the implemented dynamic systems are regarded as much greater than the simulation time step $\Delta t$.
Figure 4: Computational implementation of model dynamics. The order of the dynamics of the models used in the paper are not higher than one. Here, $x$ represents the state variable of the system. The Euler method is adopted to simulate all the dynamics. A flowchart of derivation is shown here to illustrate the approximation used in the computation.
Parameter Description Value
@@
$w_{0\mathrm{dist}}$ Intra-pool zero distance synapse strength 4
$c_{\mathrm{dist}}$ Synapse strength spatial decay coefficient 0.3
$w_{\mathrm{V1/V2b}\xrightarrow{}\mathrm{motor}}$ Uniform synapse strength from V1/V2b to motor neuron -50
$w_{\mathrm{motor}\xrightarrow{}\mathrm{V1/V2b}}$ Uniform synapse strength from motor neuron to V1/V2b 2
@@
$\mathbf{Size}_{\mathrm{pool}}$ Number of motor neurons in a motor neuron pool 2020
$v_{\mathrm{motor\_rest}}$ Motor neuron resting and reseting potential 0 mV
$v_{\mathrm{motor\_th}}$ Motor neuron threshold potential $10$ mV
$\tau_{\mathrm{motor}}$ Motor neuron membrane potential decay time constant 99 ms
$\mathrm{refractory}_{\mathrm{motor}}$ Motor neuron refractory period 55 ms
$S_{\mathrm{motor\_spike}}$ Motor neuron stochastic firing transition width $0.2$
$r_{\mathrm{Ca^{2+}}}$ $\mathrm{Ca^{2+}}]$ unit increment per action potential 11
$\tau_{\mathrm{Ca^{2+}}}$ $\mathrm{Ca^{2+}}]$ decay time constant $250$ ms
$\mathrm{Thres}_{\mathrm{Ca}^{2+}}$ $\mathrm{Ca^{2+}}]$ threshold for opening the $\mathrm{K^{+}}$ channel $10$
$S_{\mathrm{K\_chan}}$ $\mathrm{Ca^{2+}}]$ sensitivity in $\mathrm{K^{+}}$ channel $10$
$C_{\mathrm{K\_chan}}$ $\mathrm{K^{+}}$ channel induced potential drop rate $8000$ mV/s
$C_{\mathrm{limit\_position}}$ Limit position inhibitory potential drop rate $400$ mV/s
$I_{\mathrm{background},0}$ Zero-velocity background stimulation strength $1380$ mV/s
$k_{\mathrm{background},v}$ Velocity-dependent background stimulation increment coefficient $40$ mV/($\cdotm/s$)
$\mathrm{A}_{\mathrm{random}}$ Randomness amplitude of background stimulation 0.5
@@
$v_{\mathrm{V1/V2b\_rest}}$ V1/V2b resting and reseting potential 0 mV
$v_{\mathrm{V1/V2b\_th}}$ V1/V2b threshold potential $10$ mV
$\tau_{\mathrm{V1/V2b}}$ V1/V2b membrane potential decay time constant 99 ms
$\mathrm{refractory}_{\mathrm{V1/V2b}}$ V1/V2b refractory period 33 ms
$S_{\mathrm{V1/V2b\_spike}}$ V1/V2b stochastic firing transition width $0.2$
@@
$\tau_{h}$ Motor torque trace decay time constant $0.1$ s
$I_{\mathrm{thigh}}$ Torque increment for a single spike at thigh joint $0.7$ $\mathrm{N}\cdot\mathrm{m}$
$I_{\mathrm{calf}}$ Torque increment for a single spike at calf joint $1.1$ $\mathrm{N}\cdot\mathrm{m}$
$\mathrm{pos^{*}_{hip}}$ Hip joints’ target position lateral $0.1$ rad
$K_{\mathrm{P}}$ Hip controller P coefficient $30$
$K_{\mathrm{I}}$ Hip controller I coefficient $10$
@@
Parameter Description Value
@@
$w_{\mathrm{min}},w_{\mathrm{max}}]$ Weight constraint factor [0.05,0.05][-0.05,0.05]
$\tau_{\mathrm{trace}}$ Motor neuron spike trace decay time constant $0.01$ s
$\tau_{\mathrm{STDP}}$ STDP signal decay time constant 22 s
$\eta_{\mathrm{negative\_relative}}$ Relative negative learning rate in STDP signal $0.3$
$\eta$ Learning rate $5\times 10^{-10}$
$c_{\mathrm{average}}$ Averaged reward coefficient $0.5$
$\alpha_{\mathrm{vel}_{x}}$ $x$-directional velocity reward coefficient 11
$\alpha_{\omega_{\mathrm{roll}}}$ Roll angular velocity reward coefficient 0.1-0.1
$\alpha_{\omega_{\mathrm{pitch}}}$ Pitch angular velocity reward coefficient 0.1-0.1
$\alpha_{\omega_{\mathrm{yaw}}}$ Yaw angular velocity reward coefficient 0.1-0.1
@@
$r_{\mathrm{AG}}$ 2-AG production rate $1\times 10^{-3}$
$\tau_{\mathrm{AG}}$ [2-AG] decay time constant 11 s
$\mathrm{Thres\_Ca}_{\mathrm{ADO}}$ $\mathrm{Ca}^{2+}]$ threshold for adenosine release $0.3$
$r_{\mathrm{ADO}}$ Adenosine production rate $0.01$
$\tau_{\mathrm{AG}}$ Adenosine decay time constant 11 s
$\mathrm{refractory}_{\mathrm{ADO}}$ Adenosine release refractory period $0.3$ s
$\eta_{\mathrm{ADO}}$ Adenosine inhibition efficacy $1.8\times 10^{-5}$
@@
$f_{\mathrm{simulation}}$ MuJoCo simulation frequency $1000$ Hz
$L_{\mathrm{max}}$ Maximum session length $10$ s
$L_{\mathrm{non\_alive\_threshold}}$ Non-alive counter threshold $0.5$ s
$\mathrm{Thres}_{z\_\mathrm{ALIVE}}$ Alive torso vertical-ness threshold $0.5$
$\mathrm{pos^{limit,-}_{thigh,front}},\mathrm{pos^{limit,+}_{thigh,front}}$] Front thigh joint limit position $[0.6,1.4]$ rad
$\mathrm{pos^{limit,-}_{thigh,rear}},\mathrm{pos^{limit,+}_{thigh,rear}}$] Rear thigh joint limit position $[0.7,1.5]$ rad
$\mathrm{pos^{limit,-}_{calf}},\mathrm{pos^{limit,+}_{calf}}$] Calf joint limit position [1.6,1.0][-1.6,-1.0] rad
$\mathrm{Z_{torso,initial}}$ Initial torso altitude $0.35$ m
$\mathrm{frictionloss_{thigh}}$ Thigh joint frictionloss $25$
$\mathrm{frictionloss_{calf}}$ Calf joint frictionloss $10$
$\mathrm{frictionloss_{hip}}$ Hip joint frictionloss $10$